Oriole Corporation will pay dividends of $2.60, $2.95, and $3.40 in the next three years. After three years, the dividends are expected to grow at a constant rate of 3 percent per year. If the required rate of return is 13.0 percent, what is the current value of the Oriole common stock? (Round answer to 2 decimal places, e.g. 52.75.)
Answer #1
1) Calculation of stock's current price:
Year Amount PVF @13% Present value
1                       2.60 0.885                       2.30
2                       2.95 0.783                       2.31
3                       3.40 0.693                       2.36
3                     35.02 0.693                     24.27
Total                     31.24
Current price is $31.24
Working:
Calculation of dividend:
Terminal value= Dividend(1+growth)/(return-growth)
                              =3.40*(1+0.03)/(0.13-0.03)
                              = 3.502/0.10= 35.02